# Inference of multi-channel r-process element enrichment in the Milky Way using binary neutron star merger observations

### v2

@@ 4 Binary neutron star delay times inferred from Milky Way disk stars
We showed above that BNS mergers alone cannot match the Galactic stellar observations that probe Milky Way chemical abundance history, given reasonable assumptions about their astrophysical rate, population, delay time distribution and EOS. We now relax these astrophysical assumptions and ask what it would take for BNS mergers to reproduce the history of r-process chemical evolution in the Galaxy. Fig. 2 shows that the median of the predicted [Eu/Fe] abundance ratio is consistent with stellar observations around solar metallicity ([Fe/H]$\gtrsim 0.0$), but that the deviation grows with decreasing metallicity. Among the sources of uncertainty, the local merger rate, mass distribution, and neutron star EOS affect the overall abundances of r-process elements, i.e., the vertical offset in Fig. 2. On the other hand, the slope of [Eu/Fe] as a function of [Fe/H] is dominated by the merger rate evolution over time. Therefore, instead of using the delay time distribution informed by sGRBs, we try to determine the BNS DTD that could explain the observed trend of r-process abundances.
@@
Within the one-zone model connecting a BNS population to an r-process abundance history, we sample randomly in local ($z=0$) BNS rate-yield ($m_{\rm ej}R_{\mathrm{MW}}$) and DTD parameters ($t_{\rm min}$,$\alpha$) and calculate the corresponding [Eu/Fe] vs [Fe/H] track. All other population parameter uncertainties—such as those in the NS mass distribution and EOS—are implicitly marginalized over in determining the distribution of possible BNS merger rates and yields. The likelihood of the local BNS rate-yield and DTD parameters is then given by
$P(d|t_{\rm min},\alpha,m_{\rm ej}R_{\mathrm{MW}})=$
@@
×$\hskip 42.67912pt\times P({\rm[Eu/Fe]},{\rm[Fe/H]}|t_{\rm min},\alpha,m_{\rm ej}R_{\mathrm{MW}}),$
where the product is taken over all the stellar spectrum measurements and $P({\rm[Eu/Fe]},{\rm[Fe/H]}|t_{\rm min},\alpha,m_{\rm ej}R_{\mathrm{MW}})$ is a delta-function likelihood located along the r-process abundance track prediction of the model with DTD parameters ($t_{\rm min}$,$\alpha$) and local BNS rate-yield $m_{\rm ej}R_{\mathrm{MW}}$.
For this analysis, our prior is log-uniform for $t_{\rm min}\in[1,2000]$ Myr and uniform for $\alpha\in[-3.0,-0.5]$. The prior on $m_{\rm ej}R_{\mathrm{MW}}$ is informed by the rate-yield posterior from Sec. 2. In practice, we evaluate Eq. (4) via numerical integration on a grid of points drawn along the $({\rm[Eu/Fe]},{\rm[Fe/H]})$ track for each sample in $(t_{\rm min},\alpha,m_{\rm ej}R_{\mathrm{MW}})$.
@@ 5 Multi-channel r-process enrichment
We reanalyze the disk-star and SAGA observations with the sGRB-informed prior on $t_{\rm min}$ and $\alpha$, and the same prior on $m_{\rm ej}R_{\rm MW}$ as above, to constrain the second channel’s contribution. We place a uniform prior on $X_{\rm SFH}\in[0,1]$. The results are presented in Fig. 4. The disk-star observations constrain the second channel’s fractional contribution to be $X_{\rm SFH}=0.71^{+0.20}_{-0.25}$ of the integrated r-process abundance by mass. The disk-star data also tightly constrain the local BNS rate-yield to be $m_{\rm ej}R_{\mathrm{MW}}=1.9^{+2.3}_{-1.4}\;M_{\odot}/\mathrm{Myr}$. The Galactic stellar spectra do not constrain the BNS DTD relative to the sGRB prior within the multi-channel model. The good fit to the data in Fig. 4 demonstrates that a two-site model can reproduce Galactic stellar observations, whereas a sole-source BNS merger origin for r-process elements is inconsistent with the sGRB constraints on the DTD Zevin et al. (2022), as well as theoretical expectations for BNS delay times, as discussed above. The data prefer the two-channel model to the BNS-only model with a Bayes factor of $\sim 10^{5}$, computed as a Savage-Dickey density ratio of the two-channel posterior to the prior at $X_{\rm SFH}=0$, when the sGRB DTD is used as a prior. Remarkably, we find that the BNS merger channel and the star formation history-tracing channel have contributed at the same order of magnitude to the present-day r-process abundance in the Galaxy.
@@ 6 Discussion
Although BNS mergers’ rate and per event r-process element yield are consistent with astrophysical and geological measurements, a much shorter delay relative to star formation history is required for BNS mergers alone to account for the Galactic stellar [Eu/Fe] observations at different metallicities. We therefore explore the possibility of a second r-process site with no significant time delay compared to star formation. Possible candidates for such a site include collapsars (Pruet et al., 2003; Surman et al., 2006; Fujimoto et al., 2007; Siegel et al., 2019) or MHD supernovae (Nishimura et al., 2006; Winteler et al., 2012; Nishimura et al., 2017; Halevi & Mösta, 2018; Reichert et al., 2021). The second channel’s significant inferred fractional contribution of $0.71^{+0.20}_{-0.25}$ to the present-day Galactic r-process abundance is consistent with previous estimates for a collapsar channel (Siegel et al., 2019).
@@
Neutron star-black hole mergers are another possible r-process site, although their contribution is highly dependent on their rate evolution and yield (Chen et al., 2021), the latter of which may be quite low based on current evidence (Biscoveanu et al., 2023). Despite the broad range of candidates for the second channel, our study places an initial constraint on the fractional contribution from this site (or combination of sites), besides typical BNS mergers.
Systematic uncertainties in the abundance data could quantitatively change our results. The assessment of the systematic uncertainty for the data we used was done by comparison with other measurements on overlapping stars. With the 0.05–0.06 dex difference in the Eu abundance (reported as $\Delta$[Eu/H] in Battistini & Bensby (2016)), we do not expect a significant change in our results. Furthermore, the late-time abundance trend of [Eu/Fe] vs. [Fe/H] is sensitive to the power-law index of the DTD of Type Ia supernovae. However, this index of $-1.1\pm 0.1$ has been consistently inferred both from direct measurement using galaxy clusters (Freundlich & Maoz, 2021) and from measurement of Type Ia rates and the cosmic star formation history (Maoz & Graur, 2017).
@@
More detailed models for chemical evolution and star formation histories of the Milky Way will also improve the robustness of our results. In particular, they would allow for the treatment of lower-metallicity halo star observations, which are impacted by inhomogeneities in the mixing of r-process elements in the ISM. Nonetheless, we do not expect these observations to alter our qualitative conclusions; even within the one-zone model, we find similar quantitative constraints on the BNS DTD, rate-yield and second-channel contribution when the halo- and disk-star SAGA observations (Suda et al., 2011) are used in place of the Battistini & Bensby (2016) disk star ones.

### v3

@@ 4 Binary neutron star delay times inferred from Milky Way disk stars
[…]
@@
Within the one-zone model connecting a BNS population to an r-process abundance history, we sample randomly in local ($z=0$) BNS rate-yield ($m_{\rm ej}R_{\mathrm{MW}}$) and DTD parameters ($t_{\rm min}$$\alpha$) and calculate the corresponding [Eu/Fe] vs [Fe/H] track. All other population parameter uncertainties—such as those in the NS mass distribution and EOS—are implicitly marginalized over in determining the distribution of possible BNS merger rates and yields. The likelihood of the local BNS rate-yield and DTD parameters is then given by
$P(d|t_{\rm min},\alpha,m_{\rm ej}R_{\mathrm{MW}})=$ (4)
@@
$\hskip 42.67912pt\times P({\rm[Eu/Fe]},{\rm[Fe/H]}|t_{\rm min},\alpha,m_{\rm ej}R_{\mathrm{MW}}),$
where the product is taken over all the stellar spectrum measurements and $P({\rm[Eu/Fe]},{\rm[Fe/H]}|t_{\rm min},\alpha,m_{\rm ej}R_{\mathrm{MW}})$ is a delta-function likelihood located along the r-process abundance track prediction of the model with DTD parameters ($t_{\rm min}$$\alpha$) and local BNS rate-yield $m_{\rm ej}R_{\mathrm{MW}}$.
For this analysis, our prior is log-uniform for $t_{\rm min}\in[1,2000]$ Myr and uniform for $\alpha\in[-3.0,-0.5]$. The prior on $m_{\rm ej}R_{\mathrm{MW}}$ is informed by the rate-yield posterior from Sec. 2. In practice, we evaluate Eq. (LABEL:dtd_like) via numerical integration on a grid of points drawn along the $({\rm[Eu/Fe]},{\rm[Fe/H]})$ track for each sample in $(t_{\rm min},\alpha,m_{\rm ej}R_{\mathrm{MW}})$.
@@ 5 Multi-channel r-process enrichment
We reanalyze the disk-star and SAGA observations with the sGRB-informed prior on $t_{\rm min}$ and $\alpha$, and the same prior on $m_{\rm ej}R_{\rm MW}$ as above, to constrain the second channel’s contribution. We place a uniform prior on $X_{\rm SFH}\in[0,1]$. The results are presented in Fig. 4. The disk-star observations constrain the second channel’s fractional contribution to be $X_{\rm SFH}=0.71^{+0.20}_{-0.25}$ of the integrated r-process abundance by mass. In other words, 45–90% of the r-process abundance in the Milky Way today was produced by a star-formation-tracking channel. The disk-star data also tightly constrain the local BNS rate-yield to be $m_{\rm ej}R_{\mathrm{MW}}=1.9^{+2.3}_{-1.4}\;M_{\odot}/\mathrm{Myr}$. The Galactic stellar spectra do not constrain the BNS DTD relative to the sGRB prior within the multi-channel model. The good fit to the data in Fig. 4 demonstrates that a two-site model can reproduce Galactic stellar observations, whereas a sole-source BNS merger origin for r-process elements is inconsistent with the sGRB constraints on the DTD Zevin et al. (2022), as well as theoretical expectations for BNS delay times, as discussed above. The data prefer the two-channel model to the BNS-only model with a Bayes factor of $\sim 10^{5}$, computed as a Savage-Dickey density ratio of the two-channel posterior to the prior at $X_{\rm SFH}=0$, when the sGRB DTD is used as a prior. Remarkably, we find that the BNS merger channel and the star formation history-tracing channel have contributed at the same order of magnitude to the present-day r-process abundance in the Galaxy.
@@ 6 Discussion
Although BNS mergers’ rate and per event r-process element yield are consistent with astrophysical and geological measurements, a much shorter delay relative to star formation history is required for BNS mergers alone to account for the Galactic stellar [Eu/Fe] observations at different metallicities. We therefore explore the possibility of a second r-process site with no significant time delay compared to star formation. Possible candidates for such a site include collapsars (Pruet et al. (2003); Surman et al. (2006); Fujimoto et al. (2007); Siegel et al. (2019); although also see Just et al. (2022)) or MHD supernovae (Nishimura et al., 2006; Winteler et al., 2012; Nishimura et al., 2017; Halevi & Mösta, 2018; Reichert et al., 2021; Yong et al., 2021). The second channel’s significant inferred fractional contribution of $0.71^{+0.20}_{-0.25}$ (45–90%) to the present-day Galactic r-process abundance is consistent with previous estimates for a collapsar channel (Siegel et al., 2019).
@@
Neutron star-black hole mergers are another possible r-process site (Kobayashi et al., 2023), although their contribution is highly dependent on their rate evolution and yield (Chen et al., 2021), the latter of which may be quite low based on current evidence (Biscoveanu et al., 2023). Despite the broad range of candidates for the second channel, our study places an initial constraint on the fractional contribution from this site (or combination of sites), besides typical BNS mergers.
Systematic uncertainties in the abundance data could quantitatively change our results. The assessment of the systematic uncertainty for the data we used was done by comparison with other measurements on overlapping stars. With the 0.05–0.06 dex difference in the Eu abundance (reported as $\Delta$[Eu/H] in Battistini & Bensby (2016)), we do not expect a significant change in our results. Furthermore, the late-time abundance trend of [Eu/Fe] vs. [Fe/H] is sensitive to the power-law index of the DTD of Type Ia supernovae. However, this index of $1.1\pm 0.1$ has been consistently inferred both from direct measurement using galaxy clusters (Freundlich & Maoz, 2021) and from measurement of Type Ia rates and the cosmic star formation history (Maoz & Graur, 2017).
@@
More detailed models for chemical evolution and star formation histories of the Milky Way will also improve the robustness of our results. Different star formation histories could lead to horizontal shift in Figure 2, and we may underestimate the uncertainties of our inference. In particular, more detailed models would allow for the treatment of lower-metallicity halo star observations, which are impacted by inhomogeneities in the mixing of r-process elements in the ISM. Nonetheless, we do not expect these observations to alter our qualitative conclusions; even within the one-zone model, we find similar quantitative constraints on the BNS DTD, rate-yield and second-channel contribution when the halo- and disk-star SAGA observations (Suda et al., 2011) are used in place of the Battistini & Bensby (2016) disk star ones. We have repeated our inference in the one- and two-channel scenarios with a uniform star formation rate in place of the cosmic one, to give a sense of the maximum variation in the results due to the uncertainty in the star formation history. We have kept the yields per event of Fe and Eu unchanged, kept the $z=0$ rate of BNS mergers consistent with the GW constraints, and normalized the rates of CCSNe and SNe Ia to the same Fe yield per stellar mass formed relative to the scenario with the cosmic star formation history. We find that the constraints on the BNS delay-time distribution and the second-channel contribution remain largely unchanged. In particular, the second-channel contribution fraction is inferred as $0.78^{+0.15}_{-0.20}$ at 90% confidence relative to the uniform star formation history (cf. $0.71^{+0.20}_{-0.25}$ relative to the cosmic one). This is the benefit of fitting data in the abundance space [Eu/Fe] vs. [Fe/H], which singles out the properties of Eu enrichment relative to star formation, rather than the details of the absolute star formation history.
In this work, we focus on the introduction of multiple recent astrophysical constraints and the development of the Bayesian framework to properly propagate the uncertainty and quantify the contribution from different channels. In follow up work, we will use our framework to explore the impact of increasing the complexity of our modelling, by incorporating more realistic abundance yields, different star formation histories, inhomogeneous mixing of r-process elements, and variations of the DTD for both BNSs and type Ia supernovae (see, for example Kobayashi et al. (2023)).